Concurrency: Practice and Experience Papers Accepted for Publication — July 7, 1999

(C323) “Dynamic Scheduling of Process Groups”

Authors: K. Y. Wang, D. C. Marinescu, and O. F. Carbunar Current Status: Paper to Publisher (3/4/97). 34 pgs. Short Code: [Wang:97a]

(C345) “CRAFT: A Framework for F90/HPF Compiler Optimizations”

Authors: Marina Chen, James Cowie, and Jan-Jan Wu Current Status: Paper to Pubisher (4/19/99).

(C372) “Multiresolution Load Balancing in Curved Space: The Wavelet Representation” Authors: Aiichiro Nakano Current Status: Paper to Publisher (11/2/98). 20 pgs. Short Code: [Nakano:98a]

(C376) “Parallelism and Recursion in Message Passing Libraries: An Efficient Methodology” Authors: C. Rodriguez, F. de Sande, C. Leon, and L. Garcia Current Status: Paper to Publisher (11/2/980. 18 pgs. Short Code: [Rodriguez:98a]

(C380) “Parallel Execution to Embedded and Iterated Runge-Kutta Methods” Authors: Thomas Rauber and Gudula Rünger Current Status: Paper to Publisher (11/2/98). 35 pgs. Short Code: [Rauber:98a]

(C382) “COLT_{HPF}, a Coordination Layer for HPF Tasks”

Authors: Salvatore Orlando and Raffaele Perego Current Status: Paper to Publisher (2/16/99). 37 pgs. Short Code: [Orlando:98a] (C383) “Predicting the Execution Time of Message Passing Models”

Authors: J. L. Roda, C. Rodriguez, D. G. Morales, and F. Almeida Current Status: Paper to Publisher (2/16/99). 31 pgs. Short Code: [Roda:99a]

(C384) “Parallel Sensitivity Analysis for DAEs with Many Parameters”

Authors: Wenjie Zhu, and Linda Petzold Current Status: Paper to publisher (4/19/98).

(C386) “Concurrent Simulation of Neutral Flow in the GEC Reference Cell”

Authors: M. Rieffel, M. Ivanov, S. Shankar, and S. Taylor Current Status: Paper to Publisher (2/16/99). 28 pgs. Short Code: [Rieffel:98a]

(C388) “Usefulness of Adaptive Load Sharing for Parallel Processing on Networks of Workstations” Authors: Sheldon Clark and Sivarama P. Dandamudi Current Status: Paper to Publisher (2/16/99). 33 pgs. Short Code: [Clark:98a]

(C391) “Ordinary Linear Model Estimation on a Massively Parallel SIMD Computer” Authors: Erricos John Kontoghiorghes Current Status: Paper to Publisher (6/23/98). 17 pgs. Short Code: [Konto:98a]

(C397) “Reproducible Execution of SR Programs”

Authors: Ronald A. Olsson Current Status: Paper to Publisher (4/19/99).

(C400) “Distributed Control Parallelism in Multidisciplinary Aircraft Design”

Authors: Denitza T. Krasteva, Layne T. Watson, Chuck Baker, Bernard Grossman, William H. Mason, and Raphael T. Haftka Current Status: Paper to Publisher (4/19/99).

(C401) “Engineering Parallel Symbolic Programs in GpH” Authors: Hans-Wolfgang Loidl, Philip W. Trinder, Kevin Hammond, Sahalu B. Junaidu, Richard G. Morgan, Simon L. Peyton Jones Current Status: Paper to Publisher (4/19/99).

(C406) “Constructing Generic Data-Farm Templates” Authors: M. Fleury, A. C. Downton, and A. F. Clark Current Status: Paper to Pubisher (4/19/99).